OpenStudy (anonymous):

If f(x)=3x+12, find the value of the y-intercept. PLEASE HELP ME, IM BEGGING YOU.

OpenStudy (anonymous):

y=3x+12

OpenStudy (anonymous):

After that are you supposed to divide everything by 3?

OpenStudy (anonymous):

yes, divide by 3 and then you will have your y.

OpenStudy (anonymous):

Ok, so the answer I got was 12, is that correct?

Directrix (directrix):

The y-intercept is where the graph of the equation cuts the y-axis. All points on the y-axis have an x-coordinate of 0. So, y =3x+12 y = 3*0 + 12 y = 12 as you stated above. The coordinates of the y-intercept are (0, 12) @plzzhelp

OpenStudy (alekos):

also, with any straight line of the form y = mx + b b is the y-intercept and m is the slope

OpenStudy (alekos):

so no calculation required in this case

OpenStudy (anonymous):

Thank you so much!